# Safe and Efficient In-Context Learning via Risk Control

### v1

@@ 3.4 Domain-Preserving Risk Transformation
While LTT (Angelopoulos et al., 2022) supports non-monotonic losses/risks, it requires the loss to be bounded, $\ell\in[0,1]$, due to its reliance on the Hoeffding-Bentkus bound (Bentkus, 2004). Schuster et al. (2022) circumvented this by clipping all negative loss values to zero. However, for our in-context learning risk, negative losses are important: they correspond to helpful demonstrations from which we want to leverage performance gains over zero-shot. These negative losses are also quite common on our tasks, and clipping them means we lose a lot of information about the true underlying loss distribution (Fig. 4). By clipping these losses to zero, the risk control procedure cannot distinguish between performing at or better than the baseline (in our case, zero-shot). This introduces substantially more conservative early-exiting, which can make LTT impractical when we want to favor correct demonstrations for performance improvement and greater efficiency gains.
@@
    Compute $\epsilon^{\prime}=\frac{\epsilon-a}{b-a}$ and $\ell^{\prime}(\lambda;x,y,c)=\frac{\ell_{\text{ICL}}(\lambda;x,y,c)-a}{b-a}$.
@@
Figure 4: We show the distribution of our ICL loss on the TweetEval-Hate dataset with a 50% mix of correct and incorrect demonstrations. There are a significant number of negative loss values, which the loss-clipping approach sets to 0. Our risk transformation approach enables us to preserve the original underlying loss distribution.
@@
Application to In-Context Learning. Note that due to our focus on classifcation, our ICL loss is bounded $\ell_{\text{ICL}}(\lambda)\in[-1,1]$; hence we can plug in a=1,b=1a=-1,b=1 when performing loss scaling in our setting. We contrast our approach with the previous approach of clipping negative losses to zero $\ell_{\text{clip}}(\lambda):=\max\{0,\ell_{\text{ICL}}(\lambda)\}$ and empirically verify that our approach both satisfies the same risk control guarantees and achieves much greater efficiency gains (see Fig. 5 and 6).
@@ 4 Experiments
Tasks
We use a total of 8 tasks for our work, spanning three diverse domains (sentiment analysis, hate speech detection, and semantic classification). Our tasks are the Stanford Sentiment Treebank (Socher et al., 2013), FinancialPhrasebank (Malo et al., 2013), TweetEval-Hate, -Atheism, and -Feminist (Barbieri et al., 2020), AG News (Del corso et al., 2005), Text REtrieval Conference (TREC) (Li & Roth, 2002), and Unnatural (Halawi et al., 2024). A detailed description of these datasets and the domains they cover can be found in §B.
Models
We compare four models – two regular LLaMA models (Llama-3-8B (et al., 2024) and Llama-2-7B (et al., 2023)) and two LayerSkip LLaMA models (layerskip-llama3-8B and layerskip-llama2-7B (Elhoushi et al., 2024)). The LayerSkip models are additionally pre-trained to encourage the production of higher-quality intermediate representations, which provide a helpful point of comparison with the models that are not explicitly pretrained as early-exit models.
Experimental Design
In our experiments, we use the following settings:
  • In-Context Demonstrations: During the risk control calibration step, we compute a single $\hat{\lambda}$ for risk control on a 50-50 mix of both incorrect and correct demonstrations. Incorrect demonstrations are obtained by permuting the labels, as in Halawi et al. (2024).
@@ 4.1 Empirical Verification of Risk Control Guarantees
We empirically verify that our approach always controls the risk across all models and datasets, even when given a mix of correct and incorrect demonstrations in the prompts. Fig.5 shows that our approach satisfies the DFRC guarantees on risk: the in-context learning risk is controlled across all models, tasks, and risk levels $\epsilon$. We provide additional results in the appendix (§E) comparing these results with the loss-clipping approach, showing that our risk transformation approach is consistently less conservative and better matches the user-defined risk level $\epsilon$ than loss-clipping approach. This highlights both the validity of our theoretical results as well as applicability to real-world tasks. We additionally demonstrate that our risk-control guarantees hold regardless of the distribution of correct vs incorrect demonstrations in the data distribution (§G).
Figure 5: Empirical risk vs the user-specified risk level $\epsilon$ using our safe ICL model and $\ell_{\text{ICL}}$ loss over a set of mixed correct and incorrect demonstrations. Aligning with the theoretical guarantees, the risk is controlled across all models and tasks. Shaded regions correspond to one standard error over 100 experiments and are included on all plots.
@@ 4.2 Comparison of Efficiency Gains with Loss Clipping
Our adaptation of the loss-bounding approach outperforms prior LTT approaches (Jazbec et al., 2024; Schuster et al., 2022) by increasing efficiency gains while preserving the same risk control assurances. At all $\epsilon$ levels, our efficiency gains – the number of layers of computation we save by applying our approach – are strictly greater than when using clipped risk (Fig. 6), and are often significantly greater. For example, controlling the prediction gap risk with our approach at $\epsilon=0.05$ results in an average of $53\%$ less layers evaluated over all datasets and models compared to loss-clipping, while still satisfying the same rigorous risk control guarantees. Though not a primary goal of our work, it is a notable side benefit of our approach.
Figure 6: We demonstrate that our risk transformation approach enables much greater efficiency gains than the loss-clipping approach by leveraging performance gains from correct demonstrations.
4.3 Class-Conditional Effects for Correct and Incorrect Demonstrations
We examine the effects of our chosen $\hat{\lambda}$ on different sub-populations of our data, corresponding to whether the model is given correct or incorrect demonstrations. We find that where there exists a $\hat{\lambda}$ that both controls overthinking risk and preserves the accuracy gains from correct demonstrations, our approach is able to find it. In other cases, we find that there can be a direct tradeoff between taking advantage of performance gains given correct demonstrations and controlling overthinking risk given incorrect demonstrations, depending on the choice of $\hat{\lambda}$; results are shown in the appendix in Fig.16.
Though the marginal guarantees from risk control do not extend in theory to class-conditional risk control (how well our risk control works within each subgroup of our data, i.e. correct vs incorrect demonstrations) when using a mixed calibration dataset with both incorrect and correct in-context demonstrations, our approach still shows promise for finding an appropriate $\hat{\lambda}$ when one exists that can perform well for both correct and incorrect demonstrations. In practice, for applications where controlling risk given unsafe prompts is more important than performance gains from helpful prompts or achieving greater efficiency gains, we can instead use the loss-clipping approach, which is much more conservative and therefore may better control risk on unsafe prompts alone, as shown in Fig.14 and 15. However, even with loss clipping, there are no guarantees for risk control on incorrect demonstrations alone for the same reason; as shown in Fig.14 and 15, loss clipping can still violate the risk-control on the subclass of only incorrect demonstrations.
@@ 5 Related Work
In-context learning (ICL). In-context learning, also known as few-shot learning, enables LLMs to perform new tasks by conditioning on a limited number of input-output examples within the prompt without requiring gradient-based fine-tuning (Dong et al., 2024; Radford et al., 2019; Brown et al., 2020; Srivastava et al., 2023). This allows LLMs to adapt to novel tasks by generalizing from examples in the prompt, but this flexibility also makes LLMs vulnerable. For example, novice users who provide incorrect examples may make an LLM perform worse on a task than it would have without the user’s input (Halawi et al., 2024), or adversarial users can design prompts to make LLMs bypass their safeguards (Xu & Wang, 2024). In our work, we consider how the quality of in-context demonstrations can impact the safety of the model’s outputs.
Evolution of representations through LLM layers. Recent studies have begun mapping how LLM hidden representations evolve with depth, revealing structured processing phases. Cheng et al. (2025); Cheng & Antonello (2024); Cheng et al. (2023) identify a pronounced intermediate-layer “abstraction phase” in which hidden states predict brain responses to language stimuli, showing that LLMs compress inputs into low-dimensional manifolds early in processing. Layer-wise analysis has also been applied to tasks such as ciphers (Fang et al., 2025), long-context failures (Lu et al., 2024), and multilingual representations (Bafna et al., 2025; Muller et al., 2021; Wendler et al., 2024; Schut et al., 2025) in LLMs. We leverage these intermediate representations to determine when to safely make a prediction on a task.
@@
Early Exiting. Early exiting in deep neural networks enhances computational efficiency by terminating inference at intermediate layers for simpler inputs, thereby reducing resource consumption with minimal performance degradation (Teerapittayanon et al., 2016). While early exiting has been widely adopted to accelerate inference (Huang et al., 2018; Zhou et al., 2020; Elbayad et al., 2020; Han et al., 2021; Schuster et al., 2022), our work introduces a novel application of early exit architectures: mitigating the influence of incorrect demonstrations in in-context learning. Although the use of early exiting to prevent ICL overthinking has also been discussed in Halawi et al. (2024), their approach relies on static layer pruning. In contrast, our method employs dynamic, per-sample early exiting based on confidence thresholding, which enables us to make robust guarantees for safety via risk control.
@@ 6 Conclusion and Discussion
Our work introduces a novel risk-controlled early-exit framework for safe in-context learning (ICL) that robustly handles demonstrations of mixed quality. We present three important contributions to enable safety for ICL: a novel early-exit model formulation using a zero-shot baseline; a novel ICL loss designed to measure overthinking; and an adaptation of the LTT risk control framework to work for our setting. This integrated approach improves the safety, reliability, and computational efficiency of ICL under mixed-quality demonstrations.
A potential limitation of our work – and of prior work on risk control – is that we do not make conditional guarantees on risk control for correct vs incorrect demonstrations when a model is presented with demonstrations of mixed quality. Future work could investigate class-conditional risk control to provide more robust safety assurances under mixed-quality prompts. Additionally, future approaches could be applied where both correct and incorrect demonstrations are provided within the same prompt; this may reflect many real-world use cases, as a human user may have inconsistent performance even when constructing the same prompt.
Ethics Statement
This work contributes to safer deployment of LLMs by introducing a principled method to detect and mitigate “overthinking” by accounting for potential user error. Our approach additionally improves computational efficiency, making LLM inference more environmentally and economically sustainable. However, though our approach improves safety in the average case, it does not provide strong guarantees for specific subpopulations or worst-case prompts, potentially leaving some harmful scenarios unmitigated. Additionally, the increased reliance on automated risk control mechanisms may give users a false sense of security, especially if deployed without proper monitoring or human oversight.
Reproducibility Statement
Detailed descriptions of the experiment setup, including datasets, methods, and implementation details, are referenced in §4. The exact prompt formats used for all tasks are presented in §L and data processing steps are outlined in §4, with justification provided through ablation studies in §J. We provide a proof in §C showing the validity of our risk transformation approach (§3.4). Source code for all experiments can be found here: https://github.com/andreawynn/llm-risk-control
Acknowledgments
Resources used in preparing this research were provided by the Johns Hopkins + Amazon Initiative for Interactive AI, AI2AI (https://ai2ai.engineering.jhu.edu/). This work was done outside of Amazon.
Appendix A Appendix
Appendix B Detailed Description of Tasks
@@ Appendix G Risk Control at Any Proportion of Correct vs Incorrect Demos
Here, we show results demonstrating that regardless of the proportion of correct vs incorrect demos in the calibration data, we are still able to control the combined risk over a test set drawn i.i.d. from the same distribution. Results shown in Figures 10, 11 and 12 for cases when there are more correct than incorrect demonstrations (a scenario that is likely in real-world applications), but we also show in Fig.13 that even when we have many more incorrect than correct demonstrations, our risk-control guarantees still hold.

### v2

@@ 3.4 Domain-Preserving Risk Transformation
While LTT (Angelopoulos et al., 2022) supports non-monotonic losses/risks, it requires the loss to be bounded, $\ell\in[0,1]$, due to its reliance on the Hoeffding-Bentkus bound (Bentkus, 2004). (Schuster et al., 2022) circumvented this by clipping all negative loss values to zero. However, for our context-aware prediction risk, negative losses are important: they correspond to helpful context from which we want to leverage performance gains over zero-shot. These negative losses are also quite common on our tasks, and clipping them means we lose a lot of information about the true underlying loss distribution (Fig. 4). By clipping these losses to zero, the risk control procedure cannot distinguish between performing at or better than the baseline (in our case, zero-shot). This introduces substantially more conservative early-exiting, which can make LTT impractical when we want to favor helpful context for performance improvement and greater efficiency gains.
@@
    Compute $\epsilon^{\prime}=\frac{\epsilon-a}{b-a}$ and $\ell^{\prime}(\lambda;x,y,c)=\frac{\ell_{\text{c}}(\lambda;x,y,c)-a}{b-a}$.
@@
Figure 4: We show the distribution of our context-aware loss on the TweetEval-Hate dataset with a 50% mix of helpful and harmful context. There are a significant number of negative loss values, which the loss-clipping approach sets to 0. Our risk transformation approach enables us to preserve the original underlying loss distribution.
@@
Application to Context-Aware Risk Control. On all our tasks, our context-aware prediction loss is bounded by $\ell_{\text{c}}(\lambda)\in[-1,1]$; hence we can plug in a=1,b=1a=-1,b=1 when performing loss scaling in our setting. We contrast our approach with the previous approach of clipping negative losses to zero $\ell_{\text{clip}}(\lambda):=\max\{0,\ell_{\text{c}}(\lambda)\}$ and empirically verify that our approach both satisfies the same risk control guarantees and achieves much greater efficiency gains across all tasks (see Fig. 5 and 7 for in-context learning and Fig. 6 and 8 for open-ended QA).
@@ 4 Experiments
4.1 In-Context Learning
Tasks We use a total of 8 in-context learning tasks for our work, spanning three diverse domains (sentiment analysis, hate speech detection, and semantic classification). Our in-context learning tasks are the Stanford Sentiment Treebank (Socher et al., 2013), FinancialPhrasebank (Malo et al., 2013), TweetEval-Hate, -Atheism, and -Feminist (Barbieri et al., 2020), AG News (Del corso et al., 2005), Text REtrieval Conference (TREC) (Li & Roth, 2002), and Unnatural (Halawi et al., 2024). A detailed description of these datasets and the domains they cover can be found in §A.
Models We compare four models – two regular LLaMA models (Llama-3-8B (et al., 2024) and Llama-2-7B (et al., 2023)) and two LayerSkip LLaMA models (layerskip-llama3-8B and layerskip-llama2-7B (Elhoushi et al., 2024)). The LayerSkip models are additionally pre-trained to encourage the production of higher-quality intermediate representations, which provide a helpful point of comparison with the models that are not explicitly pretrained for early-exit.
We provide further details on the specifics of our experiment design for reproducibility in §B.1.
4.2 Open-Ended Language Modeling
Motivation. While our ICL experiments demonstrate risk control on classification tasks, many real-world LLM deployments involve open-ended, multi-token generation where users provide context documents. For instance, returning to our clinical scenario: a physician might upload patient notes (context) and ask the LLM to generate a treatment recommendation. If those notes contain errors or inconsistencies, we may want the LLM to recognize this and fall back to more conservative zero-shot behavior. We evaluate this setting using SQuAD v2.0, which includes both helpful context (containing the correct answer to the corresponding question) and adversarially-selected harmful context (chosen to mislead the model into giving an incorrect answer) on a question-answering (QA) task.
Task & Model. We experiment with SQuAD v2.0 (Rajpurkar et al., 2018), which includes questions with helpful context that include the correct answer to the corresponding question and questions with harmful context that are adversarially selected by human annotators to mislead the model, containing reasonable but incorrect answers. We use the T5-Large model (Raffel et al., 2020), mirroring the approach taken in CALM (Schuster et al., 2022).
We provide further details on the specifics of our experiment design for reproducibility in §B.2.
@@ 4.3 Empirical Verification of Risk Control Guarantees
Figure 5: Empirical risk vs the user-specified risk level $\epsilon$ using our safe context-aware model and $\ell_{\text{c}}$ loss over a set of mixed correct and incorrect demonstrations on our ICL tasks. Aligning with the theoretical guarantees, the risk is controlled across all models and tasks. Shaded regions correspond to one standard error over 100 experiments and are included on all plots.
Figure 6: Results of applying our risk control approach on SQuAD-v2.0 using the CALM model on different mixtures of context quality. Shaded regions correspond to one standard error over 100 experiments and are included on all plots. We demonstrate that our risk transformation approach effectively balances performance gains from helpful context with safety on misleading context while preserving robust risk-control guarantees.
We empirically verify that our approach always controls the risk across all models and datasets, even when given a mix of helpful and harmful context. Fig.5 and 6 show that our approach satisfies the DFRC guarantees on risk: the context-aware risk is controlled across all models, tasks, and risk levels $\epsilon$. We provide additional results in the appendix (§E) comparing these results with the loss-clipping approach, showing that our risk transformation approach is consistently less conservative and better matches the user-defined risk level $\epsilon$ than loss-clipping approach. This highlights both the validity of our theoretical results as well as its applicability to real-world tasks. We additionally demonstrate that our risk-control guarantees hold regardless of the distribution of helpful vs harmful context (§G).
Figure 7: We demonstrate that our risk transformation approach enables much greater efficiency gains than the loss-clipping approach by leveraging performance gains from correct demonstrations in ICL.
Figure 8: We demonstrate that our risk transformation approach enables much greater efficiency gains than the loss-clipping approach by leveraging performance gains from helpful context on the SQuAD dataset with different mixtures of context quality.
Safety of the Zero-Shot Baseline on Language Modeling Experiments. Our zero-shot fallback mechanism demonstrates selective intervention on unsafe inputs. At our risk-controlled thresholds $\hat{\lambda}$, the system defaults to zero-shot predictions 3.3x more frequently on unanswerable (misleading) questions than answerable ones (3.0% vs. 0.9%), indicating that our approach successfully discriminates on context quality. Further, when fallback occurs with unsafe (unanswerable) context, zero-shot outperforms early-exit predictions on 71% of the questions, confirming that intervention happens precisely when needed. This demonstrates a substantial safety benefit of our approach compared to CALM, which defaults to using full-model predictions that have processed the corrupted context.
Class-Conditional Effects. We examine the effects of our chosen $\hat{\lambda}$ on different sub-populations of our data, corresponding to whether the model is given helpful or harmful context. We find that where there exists a $\hat{\lambda}$ that both controls overthinking risk and preserves the accuracy gains from helpful context, our method is able to find it. In other cases, we find that there can be a direct tradeoff between taking advantage of performance gains given helpful context and controlling overthinking risk given harmful context, depending on the choice of $\hat{\lambda}$; results are shown in the appendix in Fig.20. We further describe this tradeoff in §H.
@@ 4.4 Comparison of Efficiency Gains with Loss Clipping
Our adaptation of the loss-bounding approach outperforms prior LTT approaches (Jazbec et al., 2024; Schuster et al., 2022) by increasing efficiency gains while preserving the same risk control assurances. At all $\epsilon$ levels, our efficiency gains – the number of layers of computation we save by applying our approach – are strictly greater than when using clipped risk (Fig. 7 & 8), and are often significantly greater. For example, controlling the prediction gap risk with our approach at $\epsilon=0.05$ on our in-context learning tasks results in an average of $53\%$ less layers evaluated over all datasets and models compared to loss-clipping. We achieve even more significant efficiency gains on the open-ended language modeling experiments; on SQuAD v2.0, we evaluate up to $81.8\%$ fewer layers. Though not a primary goal of our work, it is a notable benefit of our approach.
@@ 5 Related Work
Learning from Context. In-context learning (ICL), also known as few-shot learning, enables LLMs to perform new tasks by conditioning on a limited number of input-output examples within the prompt without requiring gradient-based fine-tuning (Dong et al., 2024; Radford et al., 2019; Brown et al., 2020; Srivastava et al., 2023). Methods like ICL allow LLMs to adapt to novel tasks by generalizing from context provided in the prompt. However, this flexibility also makes LLMs vulnerable, as not all context provided to a model improves performance. Novice users who provide incorrect examples of a task may make an LLM perform worse on a task than it would have without the user’s input (Halawi et al., 2024), or adversarial users can design prompts to make LLMs bypass their safeguards (Xu & Wang, 2024). In our work, we consider how the quality of in-context demonstrations can impact the safety of the model’s outputs.
Evolution of representations through LLM layers. Recent studies have begun mapping how LLM hidden representations evolve with depth, revealing structured processing phases. (Cheng et al., 2025; Cheng & Antonello, 2024; Cheng et al., 2023) identify a pronounced intermediate-layer “abstraction phase” in which hidden states predict brain responses to language stimuli, showing that LLMs compress inputs into low-dimensional manifolds early in processing. Layer-wise analysis has also been applied to tasks such as ciphers (Fang et al., 2025), long-context failures (Lu et al., 2024), and multilingual representations (Bafna et al., 2025; Muller et al., 2021; Wendler et al., 2024; Schut et al., 2025) in LLMs. We leverage these intermediate representations to determine when to safely make a prediction on a task.
@@
Early Exiting. Early exiting in deep neural networks enhances computational efficiency by terminating inference at intermediate layers for simpler inputs, thereby reducing resource consumption with minimal performance degradation (Teerapittayanon et al., 2016). While early exiting has been widely adopted to accelerate inference (Huang et al., 2018; Zhou et al., 2020; Elbayad et al., 2020; Han et al., 2021; Schuster et al., 2022), our work introduces a novel application of early exit architectures: mitigating the influence of harmful context provided to a language model. Although the use of early exiting to prevent ICL overthinking has also been discussed in (Halawi et al., 2024), their approach relies on static layer pruning. In contrast, our method employs dynamic, per-sample early exiting based on confidence thresholding, which enables us to make robust guarantees for safety via risk control.
@@ 6 Conclusion and Discussion
Our work introduces a novel risk-controlled early-exit framework for safely and robustly leveraging user-provided context of mixed quality. We present three important contributions to enable safety in this setting: a novel early-exit model formulation using a zero-shot baseline; a novel context-aware prediction loss designed to measure overthinking; and an adaptation of the LTT risk control framework to work for our setting. This integrated approach improves the safety, reliability, and computational efficiency of learning from mixed-quality context.
A potential limitation of our work – and of prior work on risk control – is that we do not make conditional guarantees on risk control for helpful vs harmful context when a model is presented with context of mixed quality. Future work could investigate class-conditional risk control to provide more robust safety assurances under mixed-quality prompts, as discussed in §H. Additionally, future approaches could be applied where both helpful and harmful context are provided within the same prompt; this may reflect many real-world use cases, as a human user may have inconsistent performance even when constructing the same prompt.
Impact Statement
This work contributes to safer deployment of LLMs by introducing a principled method to detect and mitigate “overthinking” by accounting for potentially harmful context input to LLMs. Our approach additionally improves computational efficiency, making LLM inference more environmentally and economically sustainable. However, though our approach improves safety in the average case, it does not provide strong guarantees for specific worst-case prompts, potentially leaving some harmful scenarios unmitigated. Additionally, the increased reliance on automated risk control mechanisms may give users a false sense of security, especially if deployed without proper monitoring or human oversight.
Appendix A Detailed Description of In-Context Learning Tasks
@@
Appendix B Experiment Design Details
B.1 In-Context Learning
We summarize our design as follows for the in-context learning tasks:
  1. 1.
  2. 2.
  3. 3.
    In-Context Demonstrations: During the risk control calibration step, we compute a single $\hat{\lambda}$ for risk control on a 50-50 mix of both incorrect and correct demonstrations. Incorrect demonstrations are obtained by permuting the labels, as in (Halawi et al., 2024).
  4. 4.
  5. 5.
B.2 Open-Ended Question Answering
We summarize our experiment design as follows for the SQuAD-v2.0 task:
Mirroring our in-context learning setting with incorrect and correct demonstrations, we separate the SQuAD-v2.0 questions into harmful (unanswerable) vs helpful (answerable) context. For our zero-shot model, we simply feed the questions without any context into T5-Large without any early-exiting. Since the SQuAD-v2.0 dataset does not have a list of valid answers for the unanswerable questions, we generate up to 5 answers for each of these questions with GPT-5 to measure model correctness.
During the risk control calibration step, we compute a single $\hat{\lambda}$ for risk control on a 50-50 mix of both helpful and harmful context. We use the same in-context learning risk as defined in 3.3 to evaluate our model compared to the user-defined $\epsilon$ threshold. On any example where the confidence does not exceed the $\lambda$ threshold at any layer across all tokens generated, we default to the zero-shot model for safety; we justify this strategy in §J.5.
@@ Appendix G Risk Control at Any Proportion of Correct vs Incorrect Demos
Here, we show results demonstrating that regardless of the proportion of helpful vs harmful context in the calibration data, we are still able to control the combined risk over a test set drawn i.i.d. from the same distribution. Results shown in Figures 12, 13 and 14 for cases when there are more correct than incorrect demonstrations in ICL (a scenario that is likely in real-world applications), but we also show in Fig.15 that even when we have many more incorrect than correct demonstrations, our risk-control guarantees still hold. We provide similar results on our CALM open-ended multi-token QA task in Fig. 16 and 17, once again showing that we control the risk more efficiently under varying distributions of context quality.
@@
Figure 16: Empirical risk vs the user-specified risk level $\epsilon$ using our risk transformation approach over a set of 10% answerable and 90% unanswerable context for the SQuAD-2.0 dataset.
Figure 17: Empirical risk vs the user-specified risk level $\epsilon$ using our risk transformation approach over a set of 10% answerable and 90% unanswerable context for the SQuAD-2.0 dataset.